you roll two fair dice, a white one and a green one.
find P(a number greater than 2 on the white die and exactly 4 on the green die)

Answer by stanbon(75887)

you roll two fair dice, a white one and a green one.
find P(a number greater than 2 on the white die and exactly 4 on the green die
----------------
The results on the two dice are independent.
---
P(> 2 AND 4) = P(>2)*P(4) = (4/6)*(1/6) = 4/36 = 1/9
============================
Cheers,
Stan H.
=================